# Differential Test Performance and Peer Effects

### v1

@@ 3 Empirical Results
Table 7 presents estimates and standard errors (in parentheses) for the peer effects parameter $\rho$. To provide some context and interpretation of $\rho$ consider the education production function in Equation (1). When $\rho=0$ such that there are no peer effects, it follows that the observed scores are $y_{ct}=\alpha_{c}f_{t}\mathbf{1}_{c}+X_{ct}^{c}\beta_{t}^{c}+(X_{ct}^{p}\beta_{t}^{p}+y_{ct}^{*})$ where $y_{ct}^{*}$ has the interpretation of an unobserved term that is measured in the same units as the test score variable. When $\rho$ is different from zero test score performance of student $i$ changes by $\rho$ times the leave-out-average of $y_{ct}^{*},$ through the term $\rho M_{c}y_{ct}^{*}.$ This interpretation implies that $\rho$ measures the effect on student $i$’s test score as a fraction of the increase in peer ability measured in units of test score performance. Similarly, the effect on test scores as a result of changes in peer characteristics are measured as $\rho\beta^{p}$ times the changes in relevant peer characteristics.
Standard errors are clustered at the class level. Test types of $y_{1}$ and $y_{2}$ are indicated in Columns 1 and 2. Control variables for the four specifications are outlined at the bottom of the table. Specification (1) includes only school fixed effects. Specification (2) further adds class type fixed effects. Specification (3) introduces student characteristics ($free\,lunch$, $black$, $girl$, $age$) and teacher characteristics ($black$, $master$, $experience$). Specification (4) further controls for peer characteristics, defined as the leave-out-mean of the four student characteristics variables.
@@
The comparison of the four specifications demonstrates the robustness of our estimators to additional controls. We now focus on Specification (4), which includes the full set of control variables. Across all grade levels and all test pairings, our results consistently reveal positive and statistically significant peer effects, ranging from 0.2 to 0.7. The results we find are relatively stable across the four specifications in Columns (1)-(4) with the magnitude of the coefficient $\rho$ generally decreasing somewhat for the larger models. The results for read-word are more robust in this sense especially in Grades $K$ and one. Estimated peer effects tend to be largest for specifications involving the listening score as well as the math-word score. The results for Grades $K$, two and three are quite similar, while results for Grade one are generally somewhat smaller in magnitude but remain statistically very significant. Importantly, when examining the pairing of reading and word study skills—a context wherein the tests evaluate closely related abilities as documented by our prior analysis—the effects are notably smaller than for the remaining pairings, ranging from 0.2 to 0.4 across different grade levels. The implications of these results are that improvements in peer quality measured in terms of potential SAT scores $y_{t}^{*}$ translate at the rate of 20 to 40 percent to improvements of individual scores or in other words a 10 point increase in the peer potential SAT score results in a 2 to 4 point increase in individual SAT scores. These effects are at the lower end of the spectrum of results reported in the literature.
@@
The comparison with other results in the literature is complicated by the fact that often a model with endogenous peer effects is estimated. To focus ideas, consider the case without covariates and only one outcome measure $y_{1}$. Then, an endogenous peer effect specification is $y_{1}=\lambda My_{1}+u.$ If peer effects are measured with full rather than leave-one-out means this formulation is identical to our model which is $y_{1}=\left(I+\rho M\right)u$ in this simplified stylized setting. The parameters $\rho$ and $\lambda$ satisfy the one-to-one mapping $\lambda=\rho/(1+\rho)$ which is a consequence of the fact that $\left(I+\rho M\right)=\left(I-\lambda M\right)^{-1}$. The inverse captures multiplier effects of peer performance inherent in the endogenous peer effect formulation, while $\rho$ is a summary measure of all these multiplier effects. When peer effects are measured by leave-one-out rather than full means the relationship between the two formulations is no longer exact. However, it can be shown that $\left(I+\rho M_{c}\right)^{-1}=\left(I-\lambda M_{c}\right)$ continues to hold approximately with an error that is $O\left(n_{c}^{-2}\right)$ where $n_{c}$ is the size of class $c$. Based on this approximation, we convert our measure of $\rho$ to an endogenous peer effect $\lambda=\rho/(1+\rho),$ or equivalently, $\rho=\lambda/\left(1-\lambda\right)$ for the purpose of the discussion that follows.
@@
In comparison with existing results for peer effects employing STAR data, our estimates are relatively small in magnitude but estimated with high precision and statistical significance. Using normalized kindergarten SAT scores, Graham (2008) derives endogenous peer effects of 0.46 and 0.56 for math and reading respectively, which translates to $\rho=0.86$ or $\rho=1.30$ in our context. [^12] Graham (2008) does not directly estimate endogenous peer effects. His model is $y_{ic}=\alpha_{c}+(\gamma-1)\bar{\epsilon}_{c}+\epsilon_{ic}$, with $y_{ic}$ being the outcome of student $i$ in class $c$ and $\bar{\epsilon}_{c}$ is the full mean of unobserved ability $\epsilon_{ic}$ in class $c$. Hence $\gamma-1$ is comparable with our estimate $\rho$, both are coefficients of average peer unobserved ability. Due to the full-mean specification, the equation is equivalent to $y_{ic}=\lambda\bar{y}_{c}+(1-\lambda)\alpha_{c}+\epsilon_{ic}$, where $\lambda=1-\frac{1}{\gamma}$ is the endogenous peer effects. Using normalized Kindergarten SAT math score as outcome, Rose (2017) finds an endogenous peer effect of 0.65 or a correlated effect (measured by coefficients of the peer average errors) of 0.71. When both are considered simultaneously, endogenous and correlated effects are 0.90 and -0.03, albeit with standard deviations over 1.7, indicating weak identification. Boozer and Cacciola (2001), using average percentiles of math and reading scores as outcomes and an instrumental variable methods for estimation, identify endogenous peer effects of 0.30, 0.86 and 0.92 for Grades 1 to 3 respectively, with standard deviations of 1.0, 0.12 and 0.04. Sojourner (2013) uses first-grade averages of SAT percentiles of math, reading and listening as outcome measure, and checks the exogenous peer effects of lagged peer outcomes. He finds an estimate of about 0.35 with a standard deviation around 0.14.
At the upper end of the spectrum are estimates by Lewbel, Qu, and Tang (2023) who report endogenous peer effects of 0.85 and 0.92 for small and regular classes in third-grade SAT math scores, each with standard deviations around 0.02. Translated to our parametrization, these estimates correspond to $\rho=5.6$ and $\rho=11.5$ respectively, using the approximation $\rho=\lambda/(1-\lambda)$. The magnitude of their findings is remarkably high and suggests that a mere 10-point rise in ability equivalent peer SAT scores leads to an increase of individual scores of up to over 100 points.
@@
Our proposed estimator leads to highly precise estimates compared to all the aforementioned estimates, except that of Lewbel, Qu, and Tang (2023). The enhancement in precision can be attributed in part to the elimination of individual fixed effects through differencing. Certain individual characteristics that have been recognized for their significant influence on scores are not observed in the data for Project STAR. These include parents’ education, family income and most importantly unobserved IQ. Our estimator, by removing unobserved class effects and individual fixed effects, effectively eliminates residual variance associated with these characteristics which may be one explanation for the precise estimates we obtain. In addition, our GMM estimators use additional quadratic moment restrictions efficiently.
@@
The aforementioned studies have each focused on one or two specific tests and are often limited to a particular grade. Methods that rely on random peer assignment are typically constrained in where they can be applied. In project STAR the random assignment assumption is more defensible in earlier grades (Boozer and Cacciola, 2001; Graham, 2008; Sojourner, 2013). In contrast, our analysis does not require random assignment and facilitates the estimation of peer effects across diverse test types and grade levels. Encouragingly, our findings consistently fall within a reasonably narrow range, affirming the robustness and reliability of our method.
We also provide the coefficients of peer averages of individual characteristics (age, black, free lunch, and girl) in Table 8. These four variables all pertain to test-invariant individual characteristics. The estimated coefficients can be expressed as $\rho(\beta_{v1}^{p}-f_{1}\beta_{v2}^{p})$, where $\rho\beta_{v1}^{p}$ and $\rho\beta_{v2}^{p}$ represent the exogenous peer effect of the variable on tests $t=1$ and $t=2$ respectively. Thus, these coefficients capture the differential exogenous peer effects across the two tests. We also scale the coefficients for dummies $girl$, $black$ and $free\,lunch$ by 1/100 so that the coefficients are the impact on the outcome of 1 percentage point changes in share of peers who are girl/black/free lunch receivers. The first observation is that, except for kindergarten where the presence of girls among peers positively impacts reading more than word study, none of the four variables show significant differential exogenous peer effects in the closely related word-read pair. Secondly, the share of black peers and the average peer age generally display insignificant differential exogenous peer effects. An exception is in Kindergarten, where peer age more favorably affects listening scores than reading scores. Given our general expectation of a negative impact from peers receiving free lunch and a positive impact from a higher share of girls, the results suggest that these two exogenous peer effects are stronger for reading and word study skills compared to listening and mathematical abilities in Grade 1. Similar patterns emerge in other grade levels, albeit with lower significance and consistency. Overall, the results indicate small and mostly insignificant peer effects on differential scores. This is consistent with a scenario where $y_{2}$ is essentially sufficient to control for both observed and unobserved test invariant determinants of test performance. In fact, when $f_{1}=1$ and $\beta_{v1}^{p}=\beta_{v2}^{p}$ these effects would be zero in our model and the overall evidence points to the conclusion that this is approximately the case in the sample as well.
4 Identification and Inference
4.1 Identification Conditions
We next discuss identification of our model. To establish identification we introduce additional assumptions imposed on the parameter space. We call the model identified at the parameter value $\theta_{0}$ if the population moment conditions underlying the formulation of our estimators are only satisfied at $\theta=\theta_{0}$ for all $\theta\in\Theta$. This definition can be seen as an extension of classical definitions of identification such as Rothenberg (1971) for likelihood based approaches. In general, identification is neither necessary nor sufficient for consistency which is established separately, see Newey and McFadden (1994) for a discussion, and White (1980) and Pötscher and Prucha (1997) for a discussion of the related concept of identifiable uniqueness.
@@ 4.1 Identification Conditions
We use subscript zero to denote true parameters and treat $\delta_{0}$, ignoring the restrictions given in (7), as an unrestricted parameter vector. Let $\theta_{0}=\left(\rho_{0},f_{1,0},\delta_{0}^{\prime}\right)^{\prime}$ with $\delta_{0}=\delta(f_{1,0},\rho_{0},\beta_{0})$. Then assume $\theta_{0}\in\Theta$, where $\Theta$ is a compact subset of $\Theta_{0}=\left(-1,1\right)\times(-K_{f},K_{f})\times(-K_{X},K_{X})^{p_{x}}$, and where $K_{f}$ and $K_{X}$ are finite positive constants and $p_{x}$ is the column dimension of $\underline{X}$. In addition, let $\gamma_{0}=\left(\gamma_{1,0},...,\gamma_{J,0}\right)^{\prime}$ with $\gamma_{j,0}^{2}=(\sigma_{1,0}^{2}+\sigma_{2,0}^{2}f_{1,0}^{2})\rho_{j,0}^{2}$. Then $\gamma_{0}\in\Gamma$ with $\Gamma=[K_{\gamma}^{-1},K_{\gamma}]^{J}$ where $K_{\gamma}>1$ is a finite constant.
Remark 4.1.
The assumption implies that the parameter space for $\rho$ is a subset of $[-K_{\rho},K_{\rho}]$ for some positive $K_{\rho}<1$.
In addition, we impose the the following assumption on $A$.
@@
Observe that $\Omega(\gamma)=\operatorname{diag}_{c=1}^{C}(\gamma_{\tau_{c}}^{2}I_{c})$ is diagonal for all admissible $\gamma$ and that $\Omega_{0}=\Omega(\gamma_{0})=\operatorname{Var}(u_{1}-f_{1,0}u_{2}|\mathcal{Z}_{n})$ in light of Assumptions 2 and 3 and Theorem 2.1. Also, $u^{+}(\theta_{0},\gamma)=\Omega(\gamma)^{-1/2}\Omega_{0}^{1/2}u^{+}.$ Thus it is readily seen that for all $A$ satisfying Assumption 5(i) we have
$E\left[u^{+}(\theta_{0},\gamma)|\mathcal{Z}_{n}\right]=0,$ (23)
$E\left[u^{+}(\theta_{0},\gamma)^{\prime}Au^{+}(\theta_{0},\gamma)|\mathcal{Z}_{n}\right]$ $=\operatorname{tr}\left(\Omega(\gamma)^{-1/2}A\Omega(\gamma)^{-1/2}\Omega_{0}\right)$ (24)
$=\operatorname{tr}\left(A\Omega(\gamma)^{-1}\Omega_{0}\right)$ (25)
$=0$ (26)
observing that the diagonal elements of $A\Omega(\gamma)^{-1}\Omega_{0}$ are zero.
We next discuss identification for a given matrix $H=z\equiv\left(z_{1}^{\prime},...,z_{c}^{\prime},...,z_{C}^{\prime}\right)^{\prime}$ of $\mathcal{Z}_{n}$ measurable instruments, where the $z_{c}$ are $n_{c}\times q$ matrices of random variables that are invariant at the classroom level, i.e., $z_{c}=\mathbf{1}_{c}\grave{z}_{c}$, where $\mathbf{1}_{c}$ is an $n_{c}\times 1$ vector with all elements equal to one and $\grave{z}_{c}$ is a $1\times q$ vector of class level characteristics. We focus on the just identified case with $q=1$ for ease of exposition. Additional overidentifying restrictions can be added but do not affect arguments related to identification. Consider the moment vector
$m_{n}\left(\theta,\gamma\right)=n^{-1/2}\left[\begin{array}[]{c}H^{\prime}u^{+}(\theta,\gamma)\\ u^{+}(\theta,\gamma)^{\prime}Au^{+}(\theta,\gamma)\end{array}\right]$ (27)
where $A$ is a matrix that satisfies the properties listed in Assumption 5.
@@
As a special case of (9) without covariates consider the data generating process $y_{t}=\mu^{\ast}f_{t,0}+\left(I_{n}+\rho_{0}M\right)u_{t}$ where $\mu^{*}$ is defined in (4). Suppose Assumptions 1, 2, 3, 4, 5, and 6 hold, let $m_{n}\left(\theta,\gamma\right)$ be defined in (27) and let $H=z$. In addition, assume that $q=1$ and
@@
A proof of Lemma 4.1 establishing the identifiability of $\left(\rho_{0},f_{1,0}\right)$ is given in the appendix. The subsequent corollary considers a special case of the lemma that is of interest in the context of test scores, and arises when $H=\mathbf{1}_{n}$ and when test scores are non-negative. The latter is typically the case for non-normalized test scores.
@@
Next we turn to a discussion of identification of $\theta=(\rho,f_{1},\delta^{\prime})^{\prime}$ when the model contains covariates. We use instruments $H=\left[\underline{X},z\right]$ for linear moment restrictions and a matrix $A$ for quadratic moment restrictions, with $A$ defined as before. Both $H$ and $A$ are $\mathcal{Z}_{n}$-measurable and $\operatorname{diag}(A)=0$. For ease of exposition we focus again on the case where $z$ is a vector. Recall that with covariates, $u^{+}(\theta,\gamma)$ is defined in (14).
We consider the following $(q+1)\times 1$ vector of linear and quadratic moment functions:
$m_{n}(\theta,\gamma)=\left(\begin{array}[]{c}m_{n}^{(l)}(\theta,\gamma)\\ m_{n}^{(q)}(\theta,\gamma)\end{array}\right)=n^{-1/2}\left(\begin{array}[]{c}H^{\prime}u^{+}(\theta,\gamma)\\ u^{+}(\theta,\gamma)^{\prime}Au^{+}(\theta,\gamma)\end{array}\right)$ (29)
where the functions $m_{n}^{(l)}(\theta,\gamma)$ and $m_{n}^{(q)}(\theta,\gamma)$ denote the linear and quadratic moment conditions respectively.
Note that our procedure is not designed to identify $\gamma$, nor is that required for the identification of $\theta$. Lemma 4.2 below shows that for all permissible $\gamma$ the moment condition satisfies $E\left[m_{n}(\theta_{0},\gamma)\left|\mathcal{Z}_{n}\right.\right]=0$ and $E\left[m_{n}(\theta,\gamma)\left|\mathcal{Z}_{n}\right.\right]\neq 0$ for $\theta\neq\theta_{0}$. The parameters $\gamma$ can be pinned down by estimating the variance of $u_{1}-f_{1,0}u_{2}$ for different types of groups. For identification, we further impose the following assumption.
@@
Remark.
Note that the elements of the matrix $H$ are by construction measurable w.r.t. $\mathcal{Z}_{n}=\sigma(X,z)$.
@@
Since $\underline{X}$ contains $X$, part (i) of the assumption implies that also the elements of $X$ are bounded in absolute value by $K_{H}$. Observing that $(\underline{X},z)$ has full column rank by Assumption 7(iii), as setting $\rho=0$ and $\gamma=(1,...,1)^{\prime}$ in $V(\rho,\gamma)$, (30) becomes $n^{-1}\left|E\left[z^{\prime}Q_{X}y_{2}|\mathcal{Z}_{n}\right]\right|>K_{y}>0$, where $Q_{X}=I-\underline{X}\left(\underline{X}^{\prime}\underline{X}\right){}^{-1}\underline{X}^{\prime}$. It is readily seen from an inspection of the proof of Lemma 4.2 that a necessary condition for Assumption 7(iii) is that $\left(\underline{X},E\left[\mu^{*}|\mathcal{Z}_{n}\right]\right)$ has full column rank. The latter condition is equivalent to the condition that $\left(\underline{X},E\left[y_{2}|\mathcal{Z}_{n}\right]\right)$ has full column rank, observing that in light of (9) we have $E\left[y_{2}|\mathcal{Z}_{n}\right]=E\left[\mu^{*}+\underline{X}\beta_{2}|\mathcal{Z}_{n}\right]$ for some vector of parameters $\beta_{2}.$
@@
Condition (30) generalizes condition (28) in Lemma 4.1. Under homoskedasticity or when estimators are not taking heteroskedasticity into account, the latter condition takes the more familiar form $n^{-1}\left|E\left[y_{2}^{\prime}z|\mathcal{Z}_{n}\right]\right|>K_{y}>0$. Further simplifications of (30) are obtained if in addition to heteroskedasticity, spatial correlation is ignored by setting $\rho=0.$ Considering this case is useful when discussing first step estimators for $\delta$ and $f_{1}.$ For the case with covariates (30) simplifies to $n^{-1}\left|E\left[z^{\prime}Q_{X}y_{2}|\mathcal{Z}_{n}\right]\right|>K_{y}>0$, where $Q_{X}$ is defined as above. Both versions of the simplified identification conditions are conventional rank conditions for linear instrumental variables estimators. A second scenario under which the conditions can be simplified is when $\gamma$ is evaluated at a fixed constant level. For example, when $\gamma=\bar{\gamma}$ then (28) is replaced by $n^{-1}\left|E\left[y_{2}^{\prime}\Omega(\bar{\gamma})^{-1/2}z\left|\mathcal{Z}_{n}\right.\right]\right|>K_{y}>0$ and a similar simplification obtained for (30).
The following lemma is a generalization of Lemma 4.1.
Lemma 4.2.

### v2

@@ 4.3 Peer Effects
To provide some context and interpretation of $\rho$ consider the education production function in Equation (7). When $\rho=0$ such that there are no peer effects, it follows that the observed scores are $y_{ct}=\alpha_{c}f_{t}\mathbf{1}_{c}+X_{ct}^{c}\beta_{t}^{c}+(X_{ct}^{p}\beta_{t}^{p}+y_{ct}^{*})$ where $y_{ct}^{*}$ has the interpretation of an unobserved term that is measured in the same units as the test score variable. When $\rho$ is different from zero test score performance of student $i$ changes by $\rho$ times the leave-out-average of $y_{ct}^{*},$ through the term $\rho M_{c}y_{ct}^{*}.$ This interpretation implies that $\rho$ measures the effect on student $i$’s test score as a fraction of the increase in peer ability measured in units of test score performance. Similarly, the effect on test scores as a result of changes in peer characteristics are measured as $\rho\beta^{p}$ times the changes in relevant peer characteristics.
@@
The comparison of the four specifications demonstrates the robustness of our estimators to additional controls. We now focus on Specification (4), which includes the full set of control variables. Across all grade levels, our results consistently reveal positive and statistically significant peer effects, ranging from 0.2 to 0.4. The results we find are relatively stable across the four specifications in Columns (1)-(4). The results for Grades $K$, two and three are quite similar, while results for Grade one are smaller in magnitude but remain statistically very significant. The implications of these results are that improvements in peer quality measured in terms of potential SAT scores $y_{t}^{*}$ translate at the rate of 20 to 40 percent to improvements of individual scores or in other words a 10 point increase in the peer potential SAT score results in a 2 to 4 point increase in individual SAT scores. These effects are at the lower end of the spectrum of results reported in the literature.
@@
The comparison with other results in the literature is complicated by the fact that often a model with endogenous peer effects is estimated. To focus ideas, consider the case without covariates and only one outcome measure $y_{1}$. Then, an endogenous peer effect specification for class room $c$ is $y_{ct}=\lambda M_{c}y_{ct}+u_{ct}.$ If peer effects are measured with full rather than leave-one-out means, i.e., if $M_{c}=n^{-1}\mathbf{1}_{c}\mathbf{1}_{c}^{\prime}$, this formulation is identical to our model, which is $y_{ct}=\left(I+\rho M_{c}\right)u_{ct}$ in this simplified stylized setting. The parameters $\rho$ and $\lambda$ satisfy the one-to-one mapping $\lambda=\rho/(1+\rho)$ which is a consequence of the fact that $\left(I+\rho M_{c}\right)=\left(I-\lambda M_{c}\right)^{-1}$. The inverse captures multiplier effects of peer performance inherent in the endogenous peer effect formulation, while $\rho$ is a summary measure of all these multiplier effects. When peer effects are measured by leave-one-out rather than full means the relationship between the two formulations is no longer exact. However, it can be shown that $\left(I+\rho M_{c}\right)=\left(I-\lambda M_{c}\right)^{-1}$ continues to hold approximately with an error that is $O\left(n_{c}^{-2}\right)$ where $n_{c}$ is the size of class $c$. Based on this approximation, we convert our measure of $\rho$ to an endogenous peer effect $\lambda=\rho/(1+\rho),$ or equivalently, $\rho=\lambda/\left(1-\lambda\right)$ for the purpose of the discussion that follows. Our estimates for $\rho$ in the range of $0.2$ to $0.4$ then translate to values for $\lambda$ that range from $0.16$ to $0.28$.
@@
In comparison with existing results for peer effects employing STAR data, our estimates are relatively small in magnitude but estimated with high precision and statistical significance. Using normalized kindergarten SAT scores, Graham (2008) derives endogenous peer effects of 0.46 and 0.56 for math and reading respectively, which translates to $\rho=0.86$ or $\rho=1.30$ in our context. [^13] Graham (2008) does not directly estimate endogenous peer effects. His model is $y_{ic}=\alpha_{c}+(\gamma-1)\bar{\epsilon}_{c}+\epsilon_{ic}$, with $y_{ic}$ being the outcome of student $i$ in class $c$ and $\bar{\epsilon}_{c}$ is the full mean of unobserved ability $\epsilon_{ic}$ in class $c$. Hence $\gamma-1$ is comparable with our estimate $\rho$, both are coefficients of average peer unobserved ability. Due to the full-mean specification, the equation is equivalent to $y_{ic}=\lambda\bar{y}_{c}+(1-\lambda)\alpha_{c}+\epsilon_{ic}$, where $\lambda=1-\frac{1}{\gamma}$ is the endogenous peer effects. Using normalized Kindergarten SAT math score as outcome, Rose (2017) finds an endogenous peer effect of 0.65 or a correlated effect (measured by coefficients of the peer average errors) of 0.71. When both are considered simultaneously, endogenous and correlated effects are 0.90 and -0.03, albeit with standard errors over 1.7, indicating weak identification. Boozer and Cacciola (2001), using average percentiles of math and reading scores as outcomes and an instrumental variable methods for estimation, identify endogenous peer effects of 0.30, 0.86 and 0.92 for Grades 1 to 3 respectively, with standard deviations of 1.0, 0.12 and 0.04. Sojourner (2013) uses first-grade averages of SAT percentiles of math, reading and listening as outcome measure, and checks the exogenous peer effects of lagged peer outcomes. He finds an estimate of about 0.35 with a standard deviation around 0.14.
@@
Our proposed estimator leads to highly precise estimates compared to all the aforementioned estimates, except that of Lewbel, Qu, and Tang (2023) who also report precise estimates. The enhancement in precision can be attributed in part to the elimination of individual fixed effects through differencing. Certain individual characteristics that have been recognized for their significant influence on scores are not observed in the data for Project STAR. These include parent’s education, family income and most importantly unobserved IQ. Our estimator, by removing unobserved class effects and individual fixed effects, effectively eliminates residual variance associated with these characteristics which may be one explanation for the precise estimates we obtain. In addition, our GMM estimators use additional quadratic moment restrictions efficiently.
@@
The aforementioned studies are often limited to a particular grade. Methods that rely on random peer assignment are typically constrained in where they can be applied. In project STAR the random assignment assumption is more defensible in earlier grades (Boozer and Cacciola, 2001; Graham, 2008; Sojourner, 2013). In contrast, our analysis does not require random assignment and facilitates the estimation of peer effects across grade levels. Encouragingly, our findings consistently fall within a reasonably narrow range, affirming the robustness and reliability of our method.
5 Identification
We next discuss identification of our model. To establish identification we introduce additional assumptions imposed on the parameter space. We call the model identified at the parameter value $\theta_{0}$ if the population moment conditions underlying the formulation of our estimators are only satisfied at $\theta=\theta_{0}$ for all $\theta\in\Theta$. Throughout this section we treat $f_{1}$ as an unknown parameter to be estimated. To accommodate the empirical specification in Section 3.2 and to justify the specification checks reported in Section 4.2 we establish that the parameter $f_{1}$ in (9) can be consistently estimated irrespective of whether $f_{1}=1$ holds or not. For this purpose we modify Assumption 2 to enlarge the class of models we consider to cases where $f_{1}\neq 1.$
@@ 5 Identification
Assume that Assumption 2 holds except that $E\left[y_{it}^{*}|F_{n,i,t}\right]=\kappa_{i}f_{t}$ for $t=1,2$ where $\kappa_{i}$ is measurable with respect to $\mathcal{Z}_{n}^{*}$ and $f_{t}$ is a constant.
Assumption 4 implies that the correctly specified model is (9) where we impose the normalization $f_{2}=1$ without loss of generality. When in addition $f_{1}=1$ then the correctly specified model reduces to (8). It is easy to see that the conclusions of Theorem 2.1 continue to hold under Assumption 4 when $u_{it}=y_{it}^{*}-\kappa_{i}f_{t}.$ The next assumption defines the parameter space.
Assumption 5.
We use subscript zero to denote true parameters and treat $\delta_{0}$, ignoring the restrictions given in (11), as an unrestricted parameter vector. Let $\theta_{0}=\left(\rho_{0},f_{1,0},\delta_{0}^{\prime}\right)^{\prime}$ with $\delta_{0}=\delta(f_{1,0},\rho_{0},\beta_{0})$. Then assume $\theta_{0}\in\Theta$, where $\Theta$ is a compact subset of $\Theta_{0}=\left(-1,1\right)\times(-K_{f},K_{f})\times(-K_{X},K_{X})^{p_{x}}$, and where $K_{f}$ and $K_{X}$ are finite positive constants and $p_{x}$ is the column dimension of $\underline{X}$. In addition, suppose there are $J$ class types denoted by $\tau_{c}\in\{1,\ldots,J\},$ and $\tau_{c}$ is assumed to be measurable w.r.t. $\mathcal{Z}_{n}$. Furthermore, let $\Omega(\gamma)=\operatorname{diag}_{c}^{C}\left\{\gamma_{\tau_{c}}^{2}I_{c}\right\}$ with $\gamma=(\gamma_{1},\ldots,\gamma_{J})^{\prime},$ then we assume that $\operatorname{Var}(u_{1}-f_{1,0}u_{2}|\mathcal{Z}_{n}^{*})=\Omega(\gamma_{0})$ for some $\gamma_{0}=(\gamma_{1,0},\ldots,\gamma_{J,0})^{\prime}\in\Gamma$ with $\Gamma=[K_{\gamma}^{-1},K_{\gamma}]^{J}$ where $K_{\gamma}>1$ is a finite constant.
The assumption implies that the parameter space for $\rho$ is a subset of $[-K_{\rho},K_{\rho}]$ for some positive constant $K_{\rho}<1$. The conditions postulated for $\Omega(\gamma)$ are satisfied under Assumption 9 in Appendix B regarding $u_{t}$. In addition, we impose the the following assumption on $A$.
@@
We next discuss identification for a given matrix $H=z\equiv\left(z_{1}^{\prime},...,z_{c}^{\prime},...,z_{C}^{\prime}\right)^{\prime}$ of $\mathcal{Z}_{n}$ measurable instruments, where the $z_{c}$ are $n_{c}\times q$ matrices of random variables that are invariant at the classroom level, i.e., $z_{c}=\mathbf{1}_{c}\grave{z}_{c}$, where $\mathbf{1}_{c}$ is an $n_{c}\times 1$ vector with all elements equal to one and $\grave{z}_{c}$ is a $1\times q$ vector of class level characteristics. We focus on the just identified case with $q=1$ for ease of exposition. Additional overidentifying restrictions can be added but do not affect arguments related to identification. Consider the moment vector $m_{n}\left(\theta,\gamma\right)$ defined in (19) where $A$ is a matrix that satisfies the properties listed in Assumption 6.
Observe that $\Omega(\gamma)=\operatorname{diag}_{c=1}^{C}(\gamma_{\tau_{c}}^{2}I_{c})$ is diagonal for all admissible $\gamma$ and that $\Omega_{0}=\Omega(\gamma_{0})=\operatorname{Var}(u_{1}-f_{1,0}u_{2}|\mathcal{Z}_{n})$ in light of Assumptions 4 and 5 and Theorem 2.1. Also, $u^{+}(\theta_{0},\gamma)=\Omega(\gamma)^{-1/2}\Omega_{0}^{1/2}u^{+}.$ Thus it is readily seen that for all $A$ satisfying Assumption 6(i) we have $E\left[m_{n}\left(\theta_{0},\gamma\right)\right]=0$, where the expectation of the quadratic moments is $\operatorname{tr}\left(A\Omega(\gamma)^{-1}\Omega_{0}\right)$ which is zero because diagonal elements of $A\Omega(\gamma)^{-1}\Omega_{0}$ are zero.
@@
As a special case of (13) without covariates consider the data generating process $y_{t}=\mu^{\ast}f_{t,0}+\left(I_{n}+\rho_{0}M\right)u_{t}$ where $\mu^{*}$ is defined in (4). Suppose Assumptions 3, 4, 5, 6, and 7 hold, let $m_{n}\left(\theta,\gamma\right)$ be defined in (19) and let $H=z$ where $z$ is invariant at the classroom level. In addition, assume that $q=1$ and
@@
A proof of Lemma 5.1 establishing the identifiability of $\left(\rho_{0},f_{1,0}\right)$ is given in the online appendix. The subsequent corollary considers a special case of the lemma that is of interest in the context of test scores, and arises when $H=\mathbf{1}_{n}$ and when test scores are non-negative. The latter is typically the case for non-normalized test scores.
@@
Next we turn to a discussion of identification of $\theta=(\rho,f_{1},\delta^{\prime})^{\prime}$ when the model contains covariates. We use instruments $H=\left[\underline{X},z\right]$ for linear moment restrictions and a matrix $A$ for quadratic moment restrictions, with $A$ defined as before. Both $H$ and $A$ are $\mathcal{Z}_{n}$-measurable and $\operatorname{diag}(A)=0$. For ease of exposition we focus again on the case where $z$ is a vector. Recall that with covariates, $u^{+}(\theta,\gamma)$ is defined in (18). We consider the $(q+1)\times 1$ vector of linear and quadratic moment functions $m_{n}(\theta,\gamma)$ defined in (19).
Note that our procedure is not designed to identify $\gamma$, nor is that required for the identification of $\theta$. Theorem 5.1 below shows that for all permissible $\gamma$ the moment condition satisfies $E\left[m_{n}(\theta_{0},\gamma)\left|\mathcal{Z}_{n}\right.\right]=0$ and $E\left[m_{n}(\theta,\gamma)\left|\mathcal{Z}_{n}\right.\right]\neq 0$ for $\theta\neq\theta_{0}$. The parameters $\gamma$ can be pinned down by estimating the variance of $u_{1}-f_{1,0}u_{2}$ for different types of groups. For identification, we further impose the following assumption.
@@
Note that the elements of the matrix $H$ are by construction measurable w.r.t. $\mathcal{Z}_{n}=\sigma(X,z)$. Since $\underline{X}$ contains $X$, part (i) of the assumption implies that also the elements of $X$ are bounded in absolute value by $K_{H}$. Observing that $(\underline{X},z)$ has full column rank by Assumption 8(iii), as setting $\rho=0$ and $\gamma=(1,...,1)^{\prime}$ in $V(\rho,\gamma)$, (25) becomes $n^{-1}\left|E\left[z^{\prime}Q_{X}y_{2}|\mathcal{Z}_{n}\right]\right|>K_{y}>0$, where $Q_{X}=I-\underline{X}\left(\underline{X}^{\prime}\underline{X}\right){}^{-1}\underline{X}^{\prime}$. It is readily seen from an inspection of the proof of Theorem 5.1 below that a necessary condition for Assumption 8(iii) is that $\left(\underline{X},E\left[\mu^{*}|\mathcal{Z}_{n}\right]\right)$ has full column rank. The latter condition is equivalent to the condition that $\left(\underline{X},E\left[y_{2}|\mathcal{Z}_{n}\right]\right)$ has full column rank, observing that in light of (13) we have $E\left[y_{2}|\mathcal{Z}_{n}\right]=E\left[\mu^{*}+\underline{X}\beta_{2}|\mathcal{Z}_{n}\right]$ for some vector of parameters $\beta_{2}.$
@@
Condition (25) generalizes condition (24) in Lemma 5.1. Under homoskedasticity or when estimators are not taking heteroskedasticity into account, the latter condition takes the more familiar form $n^{-1}\left|E\left[y_{2}^{\prime}z|\mathcal{Z}_{n}\right]\right|>K_{y}>0$. Further simplifications of (25) are obtained if in addition to heteroskedasticity, spatial correlation is ignored by setting $\rho=0.$ Considering this case is useful when discussing first step estimators for $\delta$ and $f_{1}.$ For the case with covariates (25) simplifies to $n^{-1}\left|E\left[z^{\prime}Q_{X}y_{2}|\mathcal{Z}_{n}\right]\right|>K_{y}>0$, where $Q_{X}$ is defined as above. Both versions of the simplified identification conditions are conventional rank conditions for linear instrumental variables estimators. A second scenario under which the conditions can be simplified is when $\gamma$ is evaluated at a fixed constant level. For example, when $\gamma=\bar{\gamma}$ then (24) is replaced by $n^{-1}\left|E\left[y_{2}^{\prime}\Omega(\bar{\gamma})^{-1/2}z\left|\mathcal{Z}_{n}\right.\right]\right|>K_{y}>0$ and a similar simplification obtained for (25). The following theorem is a generalization of Lemma 5.1.
Theorem 5.1.